OpenStudy (anonymous):

If three times a number added to 7 is divided by the number plus 8, the result is four thirds. find the number. How is this solved?

OpenStudy (anonymous):

Go backwards

OpenStudy (anonymous):

\[\frac{ 3x+7}{ x+8 }=\frac{ 4 }{ 3 }\] How would I solve this though?

OpenStudy (anonymous):

Cross multiply. 4(x + 8) = 3(3x + 7)

OpenStudy (anonymous):

Okay, got it:)Thanks!